OpenStudy (seascorpion1):

Reduction of order to find a second Frobenius series solution. Question to follow as soon as I can type it in.....

OpenStudy (seascorpion1):

\[xy''-y=0\]and \[y_1=\sum_{n=1}^{\infty}\frac{x^n}{n!(n-1)!}\]so via reduction of order:\[y_2=Cy_1\ln(x)+\sum_{n=0}^{\infty}b_nx^n\]I've substituted y2 into the ODE and got the reccurence relation:\[n(n+1)b_{n+1}-b_n=-C(\frac{2n+1}{n!(n+1)!})\]

OpenStudy (seascorpion1):

however I can't get the right substitution in the form bn=f(n)*cn (where f(n) is a term involving n and possibly C) to simplify the recurrence relation and then solve for y2. I've tried multiple substitutions and they all end up dividing by 0 meaning that my recurrence relation can not be in terms of b0 or c0. Can someone please explain the algorithm for identifying the substitution that simplifies the recurrence relation?

OpenStudy (seascorpion1):

@phi and @Directrix You've both helped me in the past, can you help me with this at all?

OpenStudy (seascorpion1):

@lochana is there any chance you could point me in the right direction with this problem? Thanks.

OpenStudy (lochana):

I am afraid. I haven't solved these kind of problems before. I need some time.:(

OpenStudy (lochana):

where does y1 come from? is it given?

OpenStudy (lochana):

https://en.wikipedia.org/wiki/Frobenius_method this might be useful

OpenStudy (seascorpion1):

y1 is not given, I calculated it. The Wikipedia link didn't help as id didn't cover reduction of order.

OpenStudy (seascorpion1):

@lochana

OpenStudy (lochana):

yes. hey, I need the original question. I am lost here:)

OpenStudy (seascorpion1):

Thanks, I'll try to send it as an attachment later.

OpenStudy (lochana):

okay. @ganeshie8 may help you with this.

OpenStudy (seascorpion1):

Thanks.

OpenStudy (seascorpion1):

Original question is attached (finally). If anyone can help with the problems I listed in my earlier posts then that would be great, thanks.

OpenStudy (seascorpion1):

@ganeshie8

OpenStudy (anonymous):

Is \(H_n\) the \(n\)th harmonic number? As in \[H_n=\sum_{k=1}^n\frac{1}{k}=1+\frac{1}{2}+\cdots+\frac{1}{n}\]

OpenStudy (seascorpion1):

Yes

OpenStudy (anonymous):

Just a quick comment for starters (I don't have enough time at the moment to look into this in much detail, but I'll come back to this question later)...\[\begin{align*}H_n-H_{n-1}&=\left(1+\frac{1}{2}+\cdots+\frac{1}{n-1}+\frac{1}{n}\right)-\left(1+\frac{1}{2}+\cdots+\frac{1}{n-1}\right)\\[1ex] &=\frac{1}{n}\end{align*}\]so you should expect the solution to the second recurrence to be given by \[b_n=\frac{1}{n(n!)(n-1)!}\]with \(b_0=1\).

OpenStudy (seascorpion1):

Thanks for getting back to me. Did you mean Hn+Hn-1 in your post above? I think I have now managed to work my way through it to the correct solution. Thanks again for your time on this. However, I do now have a new differential equation question which I'm stuck with. I've posted it in general maths (called 'ODE question to follow') and it relates to simplifying a power series substitution into an ODE. Any help would be brilliant, thanks.

OpenStudy (anonymous):

That was an unfortunate typo... But if you've figured it out, great!